# On the Structure and Semantics of Identifier Names Containing Closed Syntactic Category Words

### v1

@@ 8 Discussion and Future Work
Our analysis focuses deeply on the semantic roles of closed-category terms, including determiners, prepositions, conjunctions, and digits; as well as the behavioral codes they imply. By applying axial and selective coding to a hand-curated dataset, we uncovered patterns of usage that go beyond surface grammar and reveal how developers embed control flow, intent, and logical relationships into naming conventions.
These findings reveal that closed-category terms serve as purposeful, behaviorally expressive units of meaning that often directly map to behavioral constructs in code. Developers use these compact grammatical forms as cognitive shortcuts to signal structural distinctions, state transitions, and operational logic. This strategic use of language suggests that these terms are important to understand and study.
@@
    Closed-category terms convey rich and distinct behavioral semantics. We examined the unique roles that each type of closed-category term tends to serve. For instance, determiners often reflect selection (e.g., someItem), temporality (e.g., nextNode), or negation (e.g., noCache); conjunctions signal alternatives, conjunctional guards, or composite actions (e.g., save_and_close, key_or_iv); and digits most commonly serve to distinguish, enumerate, or encode system-specific roles (e.g., arg1, Neo4j). These behavioral distinctions have been formalized in our axial coding schema, which can serve as a foundation for recommendation systems, naming audits, or comprehension studies.
@@
    Closed-category semantics adapt to programming context in ways distinct from natural language. Prepositions that typically express motion or containment in English (e.g., to, from, on) are disproportionately found in function names, where they signal behavior, transformation, or event triggers. Conversely, determiners like no, this, and next are rare in class names but common in parameters or attributes, where positional reasoning and state scoping are essential. This contextual distribution suggests that developers adapt the semantics of closed-category terms to the structural and functional roles of identifiers, shaping how behavior is encoded in code.
@@
    Closed-category terms serve as lexical scaffolds for program comprehension. These terms often compress multiple behavioral or structural ideas into a compact name. Whether flagging control flow (e.g., ifEnabled), encoding default values (noVal), or signaling transformation (toString), they act as micro-annotations; embedding traceability, logic, or domain-specific cues directly into identifiers. This use of grammar-as-guidance offers a potentially promising direction for tools aimed at improving name clarity and developer communication.
  4. 4.
    Grammar patterns help us understand form and function. Combining axial and selective codes with grammar patterns reveals both the entities involved (open-category terms) and the behavioral relationship between them (closed-category terms). When the closed term appears at the beginning of a pattern (e.g., DT NM N, P NM N), it typically forms a unary relation, modifying a single operand (e.g., temporal status or type reinterpretation). When it appears between open terms (e.g., N P N, N CJ N), it often encodes a binary relation—linking two operands through behavior or logic (e.g., data flow or disjunction). We also note that function identifiers with a final closed term (e.g., sendTo, readFrom) may also express a binary relation, where the second operand is supplied as a parameter rather than named directly. This is consistent with observations made during data analysis in our prior work Newman_GP .
@@
Grammar patterns can also be used to scaffold naming suggestions in LLM-driven tools. Instead of generating identifiers purely from task descriptions, models could be prompted to produce names that instantiate common closed-category structures (e.g., DT NM N, P N), which are frequently associated with behavioral semantics. This approach may help align completions with human naming conventions, while still allowing flexibility in term choice.
@@ 9 Threats to Validity
Internal Validity: Abbreviations within identifiers were not expanded, which may have caused occasional misinterpretation by annotators. However, annotators had access to the surrounding source code, reducing the risk of misannotation. Grammar pattern tagging and axial coding for each closed-category term were both subject to cross-annotation by three independent annotators and evaluated using Fleiss’ Kappa to assess agreement. We used a grounded-theory approach to develop our behavioral codes. Four coders participated in open and axial coding; during the selective coding phase, one coder proposed all selective codes, which were then validated and refined collaboratively by the other three through discussion until thematic saturation was reached.
We used statistical methods to examine correlations between closed-category terms and contextual variables. We performed two chi-square tests: one to assess correlation between closed-category part-of-speech categories (e.g., Determiner, Preposition) and programming language (Java, C, C++), and another for their correlation with code context (Attribute, Function, Declaration, Parameter, Class), derived automatically via srcML collard:2016 . We applied Bonferroni correction to account for multiple comparisons. A threat to internal validity is the assumption of independence in the chi-squared test. If violated, some significance values may be distorted. However, the primary insights of RQ1, which focus on behavioral coding through qualitative analysis, are unaffected by this statistical assumption.
@@
Mitigation Strategies: To ensure transparency and reproducibility, the dataset will be made publicly available (Section 11). Annotators were allowed to inspect source code when labeling identifiers, and each identifier was independently annotated twice. Grammar patterns and axial codes were validated by multiple annotators, with inter-rater agreement assessed using Fleiss’ Kappa. We selected a representative sample from 30 software systems, sized to meet a 95% confidence level with a 5% confidence interval. Code context was derived automatically using srcML. Finally, to evaluate whether closed-category term usage varies by domain, we curated a domain-specific dataset (e.g., compilers, databases, networking tools) and compared it against a general-purpose set selected without regard to domain. We applied a Mann-Whitney U test to compare term frequencies between these groups, normalizing by lines of code to control for system size.
@@ 10 Conclusions
    A dual-level coding framework: We introduce a combined selective and axial coding scheme to interpret the semantics of determiners, prepositions, conjunctions, and digits in identifiers. This framework maps grammar structure to conceptual behavior in a way not previously formalized, and provides a solid basis for future research and development of naming practices.
@@
These findings have practical implications for code review, refactoring, and intelligent naming support. Automated tools can leverage our work to suggest context-appropriate names, detect inconsistent patterns, or provide just-in-time feedback during development. The insights we have uncovered can be packaged and curated by educators to teach semantic clarity in naming, moving beyond generic best practices to behavior-specific guidance.
@@
    Further validation of the behavioral categories we propose, especially through comprehension studies that measure how different naming structures affect developer understanding.
@@
11 Data Availability
We have created a repository that contains the data and scripts needed to generate the numbers and statistical analysis from the RQs. The scripts are in the scripts directory and need only to be run (they take no arguments). The data directory contains all of the annotation data, both broken down by closed category; each file is named after the category it contains, and amalgamated in a single file (called Tagger Open Coding). This repository can be found at this link [^5] https://github.com/SCANL/closed_category_emse_analysis_scripts.
12 Competing Interests

### v4

@@ 8 Discussion and Future Work
Closed-category terms—including determiners, prepositions, conjunctions, and numerals—are compact grammatical structures that carry dense behavioral meaning in source code. Developers use them to embed logical cues, control flow, and intent directly into identifiers. Based on our findings, we outline several actionable implications for tool builders, educators, and developers:
@@
    Treat closed-category terms as lightweight cognitive annotations. Developers should consider how they might improve identifier clarity by deliberately using closed-category terms to signal concepts such as selection (allItems), temporality (nextNode), negation (noCache), disjunction (keyOrIV), or encode system-specific roles/concepts (e.g., arg1, Neo4j). These terms encode behavioral roles that may be missed in names relying solely on open-class words. Naming checkers, LLMs, or documentation generators should highlight or recommend these roles to aid consistency and reader comprehension, as well as assist developers in reflecting on how they might improve their terminology usage in identifiers. Our axial codes offer a schema that could seed such tools.
@@
    Augment use of closed-category terms by considering code context. Our data show that closed-category term usage varies systematically across code contexts. For example, determiners like no, next, and this are very unlikely to appear in class names, while prepositions such as on or to are common in function names but rare elsewhere. As developers consider when to use closed category terms in their coding, they can leverage context to help in their decision making. Tool support can make this easier– IDEs and linters should offer context-aware naming prompts, e.g., flagging uncommon use of conjunctions in class names or suggesting appropriate determiners for declarations.
@@
    Use grammar patterns to reveal relational structure. Closed-category terms not only express behavior, they also help structure it. Patterns like DT NM N and P NM N typically signal unary relations (e.g., lastError, asFloat), while N CJ N or N P N often signal binary ones (e.g., dataOrLogger, readFromDisk). Function names with final prepositions (e.g., sendTo) sometimes encode a second operand passed via parameters. Developers should consider the latent logical structure of the grammatical patterns they use when naming. Static analyzers or naming tools could infer missing operand roles or flag relational ambiguity. For example, encountering a binary-looking pattern in a context that provides only one operand could trigger a naming prompt.
@@
Grammar patterns can also be used to scaffold naming suggestions in LLM-driven tools. Instead of generating identifiers purely from task descriptions, models could be prompted to produce names that instantiate common open- or closed-category structures (e.g., DT NM N, P N), which are frequently associated with behavioral semantics. This approach may help align completions with human naming conventions, while still allowing flexibility in term choice.
@@ 9 Threats to Validity
Internal Validity: Abbreviations within identifiers were not expanded, which may have caused occasional misinterpretation by annotators. However, annotators had access to the surrounding source code, reducing the risk of misannotation. Grammar pattern tagging and axial coding for each closed-category term were both subject to cross-validation by three independent annotators and evaluated using Fleiss’ Kappa to assess agreement. We used a grounded-theory-inspired approach to develop our behavioral codes. Four coders participated in open and axial coding; during the selective coding phase, one coder proposed all selective codes, which were then validated and refined collaboratively by the other three through discussion until thematic saturation was reached.
We used statistical methods to examine correlations between closed-category terms and contextual variables. We performed two chi-square tests: one to assess correlation between closed-category part-of-speech categories (e.g., Determiner, Preposition) and programming language (Java, C, C++), and another for their correlation with code context (Attribute, Function, Declaration, Parameter, Class), derived automatically via srcML collard:2016 . We applied Bonferroni correction to account for multiple comparisons. A threat to internal validity is the assumption of independence in the chi-squared test. If violated, some significant values may be distorted. However, the primary insights of RQ1, which focus on behavioral coding through qualitative analysis, are unaffected by this statistical assumption.
@@
Mitigation Strategies: To ensure transparency and reproducibility, the dataset will be made publicly available (Section 11.5). Annotators were allowed to inspect source code when labeling identifiers, and each identifier was independently annotated twice. Grammar patterns and axial codes were validated by multiple annotators, with inter-rater agreement assessed using Fleiss’ Kappa. We selected a representative sample from 30 software systems, sized to meet a 95% confidence level with a 5% confidence interval. Code context was derived automatically using srcML. Finally, to evaluate whether closed-category term usage varies by domain, we curated a domain-specific dataset (e.g., compilers, databases, networking tools) and compared it against a general-purpose set selected without regard to domain. We applied a Mann-Whitney U test to compare term frequencies between these groups, normalizing by lines of code to control for system size.
@@ 10 Conclusions
    A dual-level coding framework: We introduce a combined selective and axial coding scheme to interpret the semantics of determiners, prepositions, conjunctions, and numerals in identifiers. This framework maps grammar structure to conceptual behavior in a way not previously formalized, and provides a solid basis for future research and development of naming practices.
@@
These findings have practical implications for code review, refactoring, and intelligent naming support. Automated tools can leverage our work to suggest context-appropriate names, detect inconsistent patterns, or provide just-in-time feedback during the development process. The insights we have uncovered can be packaged and curated by educators to teach semantic clarity in naming, moving beyond generic best practices to behavior-specific guidance.
@@
    Further validation of the behavioral categories we propose, especially through comprehension studies that measure how different naming structures affect developer understanding. This could also extend to examining the correlation between the beahvoral categories and software quality metrics.
@@
11 Declarations
11.1 Funding
This work was not funded by any agency
11.2 Ethical approval
This work does not involve human or animal subjects and does not require IRB approval
11.3 Informed consent
This work does not involve human or animal subjects and does not require informed consent
11.4 Author Contributions
  1. 1.
    Conceptualization: Christian D. Newman
  2. 2.
    Data curation: Christian D. Newman, Anthony Peruma, Syreen Banabilah, Michael J. Decker, Reem S. AlSuhaibani, Eman Abdullah Alomar, Mahie Crabbe
  3. 3.
    Formal analysis: Christian D. Newman
  4. 4.
    Investigation: Christian D. Newman
  5. 5.
    Methodology: Christian D. Newman
  6. 6.
    Project administration: Christian D. Newman
  7. 7.
    Resources: Christian D. Newman
  8. 8.
    Software: Christian D. Newman, Farhad Akhbardeh, Anthony Peruma
  9. 9.
    Supervision: Christian D. Newman
  10. 10.
    Validation: Christian D. Newman, Anthony Peruma, Syreen Banabilah, Eman Abdullah Alomar
  11. 11.
    Visualization: Christian D. Newman
  12. 12.
    Writing – original draft: Christian D. Newman, Jonathan I Maletic, Mohamed Wiem Mkaouer, Marcos Zampieri
  13. 13.
    Writing – review & editing: Christian D. Newman, Jonathan I Maletic, Anthony Peruma, Syreen Banabilah, Michael J. Decker, Reem S. AlSuhaibani, Eman Abdullah Alomar, Marcos Zampieri
11.5 Data Availability Statement
We have created a repository that contains the data and scripts needed to generate the numbers and statistical analysis from the RQs. The scripts are in the scripts directory and need only to be run (they take no arguments). The data directory contains all annotation data broken down by closed category. Each file is named after the category it contains, and they are amalgamated in a single file (called Tagger Open Coding). This repository can be found at this link [^5] https://github.com/SCANL/closed_category_emse_analysis_scripts.
11.6 Conflict of Interests